# Deep Neural Networks and Brain Alignment: Brain Encoding and Decoding (Survey)

### v2

@@ 5.2 Metrics for Brain Decoding Models
Brain decoding methods are evaluated using popular metrics like pairwise and rank accuracy (Pereira et al., 2018; Sun et al., 2019; 2020; Oota et al., 2022c). Other metrics used for brain decoding evaluation include R2 score, mean squared error, and using Representational Similarity Matrix (Cichy et al., 2019; 2021).
@@
Pairwise accuracy. is computed as follows. The first step is to predict all the test stimulus vector representations using a trained decoder model. Let S = [S0, S1$\cdots$,Sn], $\hat{S}$ = [$\hat{S}_{0}$, $\hat{S}_{1}$$\cdots$,$\hat{S}_{n}$] denote the “true” (stimuli-derived) and predicted stimulus representations for $n$ test instances resp. Given a pair $(i,j)$ such that $0\leq i,j\leq n$, score is 1 if corr($\hat{S}_{i}$) + corr($\hat{S}_{j}$) $>$ corr(Si,$\hat{S}_{j}$) + corr(Sj,$\hat{S}_{i}$), else 0. Here, corr denotes the Pearson correlation. Figure 9 (right) illustrates the computation of Pairwise Accuracy for the case where sample $i$ and $j$ correspond to the brain activations for text stimuli “apartment” and “building” respectively. Final pairwise matching accuracy per participant is the average of scores across all pairs of test instances.
@@
Rank accuracy. is computed as follows. First, we compare each decoded vector to all the “true” stimuli-derived semantic vectors and rank them by their correlation. The classification performance reflects the rank $r$ of the stimuli-derived vector for the correct word or picture stimuli: $1-\frac{r-1}{\#instances-1}$. The final accuracy value for each participant is the average rank accuracy across all instances.
@@ 6 Brain Encoding
Encoding is the learning of the mapping from the stimulus domain to the neural activation. The quest in brain encoding is for “reverse engineering” the algorithms that the brain uses for sensation, perception, and higher-level cognition. The foundational approach to constructing a brain encoder, illustrated in Figure 11, adopts a general brain alignment strategy previously implemented in several notable studies (Jain & Huth, 2018; Toneva & Wehbe, 2019; Aw & Toneva, 2023; Oota et al., 2024b). This method predicts fMRI recordings at every voxel for each participant, utilizing DNN representations that mirror the participant’s engagement in tasks such as reading or listening.
Building on this foundation, the recent advancements in neuroimaging technologies have enhanced our ability to closely approximate how the brain responds to different stimuli, thereby deepening our understanding of the brain’s information processing mechanisms. Concurrently, advancements in deep neural network (DNN) models have led to the development of highly efficient models across different modalities, including language, vision, speech, and multimodal interactions. These models have set new benchmarks in performance for a wide range of applications. Leveraging cutting-edge neuroimaging techniques and DNN models, this section offers a comprehensive review of the task settings for brain encoding and thelatest achievements in understanding language processing, visual object recognition, auditory perception, and multimodal processing in the brain.
In the discussion on encoding task settings, we present stimulus downsampling, TR alignment, and voxelwise encoding models. In linguistic brain encoding, we explore recent breakthroughs in applied Natural Language Processing (NLP) that facilitate the reverse engineering of the language function of the brain. In the realm of vision brain encoding, pioneering results have been achieved in reverse engineering the function of the ventral visual stream for object recognition, thanks to the advancements and impressive successes of deep Convolutional Neural Networks (CNNs) and Vision Transformers. Additionally, we present the latest insights into auditory and multimodal brain encoding. This systematic approach informs the organization of this section. Overall, Figure 10 classifies the encoding literature along various stimulus domains such as vision, auditory, multimodal, and language and the corresponding tasks in each domain. Finally, Table 3 summarizes various encoding models proposed in the literature related to textual, audio, visual, and multimodal stimuli.
@@
the Syntax and Semantics, level2, l sep=10mm,[ (Wang et al., 2020; Zhang et al., 2022a; Caucheteux et al., 2021; Toneva et al., 2022a; Reddy & Wehbe, 2021; Toneva et al., 2022b; Oota et al., 2023a), tnode] ] [NLP Tasks in
Language Models and Brains, level2, l sep=10mm, [ (Gauthier & Levy, 2019; Schwartz et al., 2019; Toneva et al., 2020; Oota et al., 2022b; Sun & Moens, 2023; Sun et al., 2023; Loong Aw et al., 2023; Kumar et al., 2022; Aw & Toneva, 2023; Merlin & Toneva, 2022; Oota et al., 2024b; Tuckute et al., 2024; Kauf et al., 2024b), tnode] ] [Alignment Between Pretrained Language
Models and Brain , level2, l sep=10mm, [ (Huth et al., 2016; Anderson et al., 2017a; Jain & Huth, 2018; Toneva & Wehbe, 2019; Deniz et al., 2019; Schrimpf et al., 2021; Caucheteux & King, 2020; Goldstein et al., 2022; Antonello et al., 2021; Oota et al., 2022a; AlKhamissi et al., 2024), tnode] ] ] [Multimodal, level1, l sep=10mm, [Incorporating
Language into Vision, level2, l sep=10mm,[ (Oota et al., 2022e; Popham et al., 2021; Tang et al., 2024; Wang et al., 2022), tnode] ] [Language & Auditory and Vision, level2, l sep=10mm,[ (Lu et al., 2022; Dong & Toneva, 2023; Nakagi et al., 2024), tnode] ] ] [Auditory, level1, l sep=10mm, [Alignment
Between Pretrained Speech Models and Brains, level2, l sep=10mm,[ (Vaidya et al., 2022; Millet et al., 2022; Tuckute et al., 2023; Oota et al., 2023c), tnode] ] [Speech Tasks and Brains, level2, l sep=10mm,[ (Tuckute et al., 2023; Oota et al., 2023d; Li et al., 2024), tnode] ] ] [Vision, level1, l sep=10mm, [Alignment Between
Vision Models and Brains , level2, l sep=10mm,[ (Yamins et al., 2014; Eickenberg et al., 2017; Cichy et al., 2016; Kubilius et al., 2019; Conwell et al., 2023; Choi et al., 2024; Wasserman et al., 2024), tnode] ] [Vision Tasks and Brains, level2, l sep=10mm,[ (Wang et al., 2019; Dwivedi et al., 2021; Khosla & Wehbe, 2022), tnode] ] ] ]
Figure 10: Categorization of Brain Encoding Studies
Figure 11: Scheme for Brain Encoding (top): this approach learns a function to predict the fMRI recordings at every voxel of each participant using the model representations that correspond to the same text read or listened by the participant. Ridge regression vs. Banded ridge regression (bottom) plot was adapted from la Tour et al. (2022). Each color (or band) represents a different feature space.
@@ 6.1 Encoding Task Settings
Stimulus downsampling. In the context of narrative story reading or listening, the rate of fMRI data acquisition was lower than the rate at which the text stimulus was presented to the subjects, several words fall under the same TR in a single acquisition. Hence, previous studies match the stimulus acquisition rate to fMRI data recording by downsampling the stimulus features using a 3-lobed Lanczos filter (Huth et al., 2016; Jain & Huth, 2018; Toneva & Wehbe, 2019; Antonello et al., 2021; Oota et al., 2024b). After downsampling, word-embeddings corresponding to each TR are obtained.
@@
For the naturalistic audio, (Vaidya et al., 2022; Antonello et al., 2024) windowed the stimulus waveform with a sliding window of size 16 s and stride 100ms before feeding it into the model. Further, the features are downsampled as previously described, using Lanczos interpolation, to match with sampling rate of fMRI recordings.
@@
Similarly for the naturalistic videos, the rate of fMRI data acquisition (TR = 2 seconds) in the shortclips dataset (Huth et al., 2022) is lower than the rate at which the stimulus was presented to the subjects (15 frames per second), 30 frames of a video were viewed under the same TR for a single fMRI acquisition (Popham et al., 2021). This helps synchronization between the stimulus presentation rate and fMRI data recording, which we then leverage to train our encoding models.
@@
fMRI Time Repetition (TR) alignment. To account for the slowness of the hemodynamic response, in general, previous studies model the HRF using a finite response filter (FIR) per voxel and for each subject separately with a delay of 8 to 12 secs (Jain & Huth, 2018; Toneva & Wehbe, 2019; Popham et al., 2021; Oota et al., 2024b; Antonello et al., 2024). Table 2 summarizes current brain encoding studies with a fixed HRF delay.
@@
MEG preprocessing and alignment. The minimal processing steps described in Gwilliams et al. (2023) are as follows. On raw MEG data and for each subject separately, using MNE-Python [^5] https://mne.tools/stable/index.html defaults parameters, the following steps should be executed:
  • bandpass filtered the MEG data between 0.5 and 30.0 Hz,
  • temporally-decimated the data 10x
  • segmented these continuous signals between -200 ms and 600 ms after word onset (note: this continuous signals varies for phoneme onset)
  • applied a baseline correction between -200 ms and 0 ms, and
@@
  • clipped the MEG data between fifth and ninety-fifth percentile of the data across channels.
In contrast to the fMRI recordings, MEG recordings have much higher time resolution. Epoching and downsampling MEG data can result in aligned word-level or phoneme-level brain recordings (Gwilliams et al., 2023; Toneva et al., 2020; Oota et al., 2023b).
Figure 12: Variance partitioning approach (Lescroart, 2017) (left) and Residual approach (Toneva et al., 2022a) (right).

### v3

@@ 5.2 Metrics for Brain Decoding Models
Brain decoding methods are evaluated using several popular metrics, including pairwise accuracy and rank accuracy (Pereira et al., 2018; Sun et al., 2019; 2020; Oota et al., 2022c). Additionally, other metrics such as R2 score, mean squared error, and Representational Similarity Matrix are also employed (Cichy et al., 2019; 2021).
Pairwise accuracy. This metric is computed as follows:
  1. 1.
    Predict all test stimulus vector representations using a trained decoder model.
  2. 2.
    Let $\text{X}=[\text{X}_{0},\text{X}_{1},\cdots,\text{X}_{n}]$ and $\hat{\text{X}}=[\hat{\text{X}}_{0},\hat{\text{X}}_{1},\cdots,\hat{\text{X}}_{n}]$ denote the “true” (stimuli-derived) and predicted stimulus representations for $n$ test instances, respectively.
  3. 3.
    For a given pair $(i,j)$ where $0\leq i,j\leq n$, the score is 1 if:
    $\text{corr}(\text{X}_{i},\hat{\text{X}}_{i})+\text{corr}(\text{X}_{j},\hat{\text{X}}_{j})>\text{corr}(\text{X}_{i},\hat{\text{X}}_{j})+\text{corr}(\text{X}_{j},\hat{\text{X}}_{i})$ (5)
    Otherwise, the score is 0. Here, corr denotes the Pearson correlation.
@@
Figure 8 (right) illustrates the computation of Pairwise Accuracy for the case where samples $i$ and $j$ correspond to the brain activations for text stimuli “apartment” and “building” respectively. The final pairwise matching accuracy per participant is the average of scores across all pairs of test instances.
Rank accuracy. This metric is computed as follows:
  1. 1.
@@
    Compare each decoded vector to all the “true” stimuli-derived semantic vectors and rank them by their correlation.
  2. 2.
    The classification performance reflects the rank $r$ of the stimuli-derived vector for the correct word or picture stimuli:
    Rank $\text{Rank accuracy}=1-\frac{r-1}{\text{\#instances}-1}$ (6)
  3. 3.
    The final accuracy value for each participant is the average rank accuracy across all instances.
These metrics provide comprehensive evaluation of brain decoding models, assessing their ability to accurately reconstruct stimuli from brain activity patterns.
5.3 Representational Similarity Metrics for Brain Encoding and Decoding Models
Three popular representational similarity metrics commonly used in brain encoding and decoding studies (Anderson et al., 2017b; Marques et al., 2021; AlKhamissi et al., 2024) are representational dissimilarity matrices (RDM) (Kriegeskorte et al., 2008), centered kernel alignment (CKA) (Kornblith et al., 2019) and normalized distribution similarity (NDS). Unlike brain encoding or decoding models that rely on linear regression, both of these metrics are nonparametric and do not require any additional training.
[…]
  1. 1.
    To construct the RDM, compute the pairwise dissimilarities between all pairs of responses for the $M$ different stimuli.
  2. 2.
[…]
    $\text{D}_{ij}=1-\frac{\text{cov}(\vec{r}_{i},\vec{r}_{j})}{\sqrt{\text{var}(\vec{r}_{i})\cdot\text{var}(\vec{r}_{j})}}$ (7)
  3. 3.
[…]
    $\text{B}_{ij}=1-\frac{\text{cov}(\vec{b}_{i},\vec{b}_{j})}{\sqrt{\text{var}(\vec{b}_{i})\cdot\text{var}(\vec{b}_{j})}}$ (8)
  4. 4.
    Finally, RSA involves calculating the correlation between the two RDMs D and B. This correlation quantifies the similarity between the representational spaces of the two systems (brain activity and the model).
Centered Kernel Alignment (CKA).  Kornblith et al. (2019) introduced CKA as a substitute for Canonical Correlation Analysis (CCA) to assess the similarity between neural network representations. Unlike linear predictivity, it is a nonparameteric metric and therefore does not require any additional training. CKA is particularly effective with high-dimensional representations, and is reliabile in identifying correspondences between representations in networks trained from different initializations (Kornblith et al., 2019). The linear CKA metric is computed as follows:
  1. 1.
    CKA between input matrices X and Y is computed as
    $\text{CKA}(\text{X},\text{Y})=\frac{\text{HSIC}(\text{K},\text{L})}{\sqrt{\text{HSIC}(\text{K},\text{K})\text{HSIC}(\text{L},\text{L})}}$ (9)
    where K and L are the similarity matrices computed as $\text{K}=\text{X}\text{X}^{\text{T}}$ and $\text{L}=\text{Y}\text{Y}^{\text{T}}$, respectively.
  2. 2.
[…]
  3. 3.
    CKA values range from 0 to 1, where CKA = 1 indicates perfect similarity, and CKA = 0 indicates no similarity.
Normalized Distribution Similarity (NDS).  Marques et al. (2021) introduced normalized distribution similarity. The normalized distribution similarity score is calculated as follows:
$\text{Score}=\frac{\text{D}_{m}}{\text{D}_{b}}\$ (10)
$\text{D}_{m}=1-\text{KS}_{c}^{\text{M}-\text{E}}$
$\text{D}_{b}=1-\text{KS}_{c}^{\text{E}-\text{E}}$
where $\text{D}_{m}$ is the distance Kolmogorov-Smirnov (KS) between the empirical model distribution (M) and the empirical biological distribution (E), and $\text{D}_{b}$ is an estimate of the expected distance from KS and c denotes the ceiling. A low score indicates that the model distribution does not match the biological distribution, while a score of 1 means that the model distribution is indistinguishable from the biological distribution when considering experimental variability.
5.4 Noise Ceiling Metric for Quantifying Model Performance
In linguistic or auditory or vision brain encoding studies, researchers aim to better interpret the accuracy of model-level predictions by quantifying the maximal possible prediction performance. This is often achieved by treating inter-participant variability as a “noise ceiling” representing the portion of neural response variability that cannot be predicted by a computational model without access to individual participant measurements (Schrimpf et al., 2020; 2021; Tuckute et al., 2024; Oota et al., 2024a; AlKhamissi et al., 2024).
Several studies compute the noise ceiling by using repeated trials of stimuli to obtain the maximum explained variance (Schoppe et al., 2016; Deniz et al., 2019; Popham et al., 2021). By presenting the same stimulus multiple times and recording the neural responses, researchers can estimate the consistency of responses within and between participants. This approach helps distinguish the variance attributable to the stimulus from the variance due to noise or individual differences. Incorporating noise ceiling into evaluation metrics provides a benchmark for interpreting model performance, allowing for a more accurate assessment of how much of the neural data the model can potentially explain.
This consideration is crucial when comparing models to neural data because it sets an upper limit on the expected performance. Without accounting for the noise ceiling, a model’s predictive accuracy might be underestimated due to the inherent variability in neural responses that no model could capture. By normalizing model performance relative to the noise ceiling, researchers can more fairly compare different models and better understand the extent to which a model approximates the true neural processing of linguistic, auditory or visual information.
@@ 6 Brain Encoding
Encoding is the process of learning the mapping from the stimulus domain to neural activation. The quest in brain encoding is for “reverse engineering” the algorithms that the brain uses for sensation, perception, and higher-level cognition. The foundational approach to constructing a brain encoder, illustrated in Figure 10(a), adopts a general brain alignment strategy previously implemented in several notable studies (Jain & Huth, 2018; Toneva & Wehbe, 2019; Aw & Toneva, 2023; Oota et al., 2024b). This method predicts brain recordings at every voxel for each participant, utilizing DNN representations that mirror the participant’s engagement in tasks such as reading or listening.
Building on this foundation, recent advancements in neuroimaging technologies have enhanced our ability to closely approximate how the brain responds to different stimuli, thereby deepening our understanding of the brain’s information processing mechanisms. Concurrently, advancements in deep neural network (DNN) models have led to the development of highly efficient models across different modalities, including language, vision, speech, and multimodal interactions. These models have set new benchmarks in performance for a wide range of applications. Leveraging cutting-edge neuroimaging techniques and DNN models, this section offers a comprehensive review of the task settings for brain encoding and the latest achievements in understanding language processing, visual object recognition, auditory perception, and multimodal processing in the brain.
In the discussion on encoding task settings, we present stimulus downsampling, TR (Repetition Time) alignment, and voxelwise encoding models. In linguistic brain encoding, we explore recent breakthroughs in applied Natural Language Processing (NLP) that facilitate the reverse engineering of the language function of the brain. In the realm of vision brain encoding, pioneering results have been achieved in reverse engineering the function of the ventral visual stream for object recognition, thanks to the advancements and impressive successes of deep Convolutional Neural Networks (CNNs) and Vision Transformers. Additionally, we present the latest insights into auditory and multimodal brain encoding. This systematic approach informs the organization of this section. Overall, Figure 9 classifies the encoding literature along various stimulus domains such as vision, auditory, multimodal, and language and the corresponding tasks in each domain. Finally, Table 6 summarizes various encoding models proposed in the literature related to textual, audio, visual, and multimodal stimuli.
@@
the Syntax and Semantics, level2, l sep=10mm,[ (Wang et al., 2020; Zhang et al., 2022b; Caucheteux et al., 2021a; Toneva et al., 2022a; Reddy & Wehbe, 2021; Toneva et al., 2022b; Oota et al., 2023a), tnode] ] [NLP Tasks in Language Models and Brains, level2, l sep=10mm, [ (Gauthier & Levy, 2019; Schwartz et al., 2019; Toneva et al., 2020; Oota et al., 2022b; Sun & Moens, 2023; Sun et al., 2023; Loong Aw et al., 2023; Kumar et al., 2022; Aw & Toneva, 2023; Merlin & Toneva, 2022; Oota et al., 2024b; Tuckute et al., 2024; Kauf et al., 2024b), tnode] ] [Pretrained Language
Models and Brain , level2, l sep=10mm, [ (Huth et al., 2016; Anderson et al., 2017a; Jain & Huth, 2018; Toneva & Wehbe, 2019; Deniz et al., 2019; Schrimpf et al., 2021; Caucheteux & King, 2022; Goldstein et al., 2022; Antonello et al., 2021; Oota et al., 2022a; AlKhamissi et al., 2024), tnode] ] ] [Multimodal, level1, l sep=10mm, [Multimodal models and Unimodal Stimuli, level2, l sep=10mm,[ (Oota et al., 2022e; Popham et al., 2021; Tang et al., 2024; Wang et al., 2023), tnode] ] [Multimodal models and Multimodal Stimuli, level2, l sep=10mm,[ (Lu et al., 2022; Dong & Toneva, 2023; Nakagi et al., 2024), tnode] ] ] [Auditory, level1, l sep=10mm, [Pretrained Speech Models and Brains, level2, l sep=10mm,[ (Vaidya et al., 2022; Millet et al., 2022; Tuckute et al., 2023; Oota et al., 2023c), tnode] ] [Speech Tasks and Brains, level2, l sep=10mm,[ (Tuckute et al., 2023; Oota et al., 2023d; Li et al., 2024b), tnode] ] ] [Vision, level1, l sep=10mm, [Vision Models and Brains , level2, l sep=10mm,[ (Yamins et al., 2014; Eickenberg et al., 2017; Cichy et al., 2016; Kubilius et al., 2019; Conwell et al., 2024; Choi et al., 2024; Wasserman et al., 2024), tnode] ] [Vision Tasks and Brains, level2, l sep=10mm,[ (Wang et al., 2019; Dwivedi et al., 2021; Khosla & Wehbe, 2022), tnode] ] ] ]
Figure 9: Categorization of Brain Encoding Studies Across Different Modalities and Tasks. The categorization is organized by stimulus modalities (e.g., language, vision, auditory, and multimodal) and the nature of the representations used—either from pretrained models or from task-specific models.
(a) Encoding schema
(b) Ridge regression vs. Banded ridge regression
Figure 10: (a) Scheme for Brain Encoding (top): This approach learns a function to predict the brain recordings at every voxel of each participant using the model representations that correspond to the same text read or listened by the participant. (b) Ridge regression vs. Banded ridge regression (bottom) plot is adapted from (la Tour et al., 2022), and used with permission from the respective authors. Each color (or band) represents a different feature space.
@@ 6.1 Encoding Task Settings
Stimulus downsampling. In the context of narrative story reading or listening, the rate of fMRI data acquisition is lower than the rate at which the text stimulus is presented to the subjects; several words fall under the same TR in a single acquisition. Therefore, previous studies match the stimulus acquisition rate to fMRI data recording by downsampling the stimulus features using a 3-lobed Lanczos filter (Huth et al., 2016; Jain & Huth, 2018; Toneva & Wehbe, 2019; Antonello et al., 2021; Oota et al., 2024b). After downsampling, word embeddings corresponding to each TR are obtained.
@@
For naturalistic audio, Vaidya et al. (2022); Antonello et al. (2024b) windowed the stimulus waveform with a sliding window of size 16s and stride 100 ms before feeding it into the model. Further, the features are downsampled as previously described, using Lanczos interpolation, to match the sampling rate of fMRI recordings.
@@
Similarly, for naturalistic videos, the rate of fMRI data acquisition (TR = 2 seconds) in the shortclips dataset (Huth et al., 2022) is lower than the rate at which the stimulus was presented to the subjects (15 frames per second). Consequently, 30 frames of a video were viewed during the same TR for a single fMRI acquisition (Popham et al., 2021). This synchronization between the stimulus presentation rate and fMRI data recording is then leveraged to train our encoding models.
fMRI preprocessing. The minimal preprocessing steps described in fMRIPrep framework (Esteban et al., 2019) to obtain the preprocessed fMRI data from raw BOLD fMRI recordings are as follows. For raw fMRI data and for each subject separately, using fMRIPrep, the following steps should be performed:
  • Take the raw functional MRI data in either DICOM or NIfTI format, and ensure that all DICOM or NIfTI files from the fMRI scan are organized in a single folder. DICOM files typically consist of a series of 2D images for each slice and time point, whereas a NIfTI file is a 4D file containing all the volumes (time points) of the scan. It is optional to start directly with DICOM files or first convert DICOM files to NIfTI format for preprocessing.
  • BOLD reference image estimation:
    1. 1.
      Mean Functional Image: The mean of all functional volumes is computed to create a stable and representative reference image.
    2. 2.
      Single Functional Volume: Alternatively, a single high-quality volume (e.g., the first or a middle time point) may be selected as the reference.
  • Motion Correction (Realignment): All functional images are aligned to the reference image that correct for translational and rotational movements.
  • Coregistration: The reference functional image is aligned with the participant’s anatomical (structural) MRI scan. This step bridges functional and anatomical data, allowing for accurate localization of brain activity.
  • Normalization: Images are transformed into a standardized space (e.g., MNI space) to allow for group analyses across participants.
  • Smoothing (Optional): Apply a Gaussian kernel to smooth the data, enhancing the signal-to-noise ratio (SNR).
  • Confound regression: Perform confound regression on both smoothed and unsmoothed data to mitigate artifacts from head motion and physiological noise.
@@
fMRI Repetition Time (TR) alignment. To account for the slowness of the hemodynamic response, previous studies generally model the Hemodynamic Response Function (HRF) using a finite impulse response (FIR) filter per voxel and for each subject separately, with a delay of 8 to 12 seconds (Jain & Huth, 2018; Toneva & Wehbe, 2019; Popham et al., 2021; Oota et al., 2024b; Antonello et al., 2024b). Table 5 summarizes current brain encoding studies with a fixed HRF delay.
@@
MEG preprocessing and alignment. Several encoding and decoding studies using MEG recordings Toneva et al. (2020; 2022a); Gwilliams et al. (2023); Hebart et al. (2023); Benchetrit et al. (2024) describe the following minimal preprocessing steps. For raw MEG data and for each subject separately, using MNE-Python library (Gramfort et al., 2013) [^5] https://mne.tools/stable/index.html, the following steps should be executed:
  • Bandpass filtering in MEG analysis involves allowing only a specific range of frequencies to pass through, filtering out lower frequencies (such as slow drifts) and higher frequencies (like noise or muscle artifacts). Specifically, in language ERP (Event-Related Potential) studies (Widmann & Schröger, 2012; Tanner et al., 2015), using a high-pass filter above 0.3 Hz introduced significant effects too early compared to unfiltered data. Based on this, these studies suggest an optimal high-pass value of 0.1 Hz for language processing. Therefore, the choice of frequency range for bandpass filters depends on the brain signals being studied, with prior studies typically using a range between 0.1 and 40 Hz (Toneva et al., 2020; Gwilliams et al., 2023; Hebart et al., 2023).
  • Temporally decimating the data or downsampling it over time reduces the computational load for downstream analyses. For example, if MEG data is originally sampled at 1,000 Hz (1,000 data points per second) and is decimated by a factor of 10, the resulting data will have a sampling rate of 100 Hz (100 data points per second).
  • Segmenting the continuous signals into smaller, distinct time intervals or epochs is often done to isolate specific events or time windows of interest for further analysis. For instance, Toneva et al. (2022a) analyzed the data from the beginning of the stimulus word presentation (i.e., 0 ms) to 800 ms, while Hebart et al. (2023) used the signal from an optical sensor to epoch the continuous data from –100 ms to 1,300 ms relative to stimulus onset.
  • Applying baseline correction adjusts the signal relative to the pre-stimulus period, removing any potential drift or noise.
@@
  • Clipping the MEG data between the fifth and ninety-fifth percentiles across channels reduces the impact of extreme values or outliers.
Figure 11: Methods for Interpreting Neural Representations in Brain Encoding Models. This figure illustrates two key approaches used to analyze the contribution of different features in neural encoding: (Left) Variance partitioning approach (Lescroart, 2017) and (Right) Residual approach (Toneva et al., 2022a). Adapted from (Toneva et al., 2022a), and used with permission from the respective authors.